OpenStudy (anonymous):

How would I find the Vertical and Horizontal Asymptotes of the following, using Limits?

OpenStudy (anonymous):

\(\ \Huge y=\frac{x}{x-2}. \)

OpenStudy (anonymous):

Also, How would I know whether to approach the HA from above or below for \(\ x=\pm\infty \)?

OpenStudy (anonymous):

When Graphing

OpenStudy (anonymous):

horizontal asymptotes are defined as the limit of a function, L, as x approaches inf and negative inf.

OpenStudy (anonymous):

you'll have to look at both postive infinity and negative infinity... you might have two horizontal asymptotes..

OpenStudy (anonymous):

Also, I'm a bit confused on finding the x/y intercepts.... Where do you plug in zero?

OpenStudy (anonymous):

for vertical asymptotes, it's just the "opposite".... that is, take the limit of x/(x-2) as x approaches 2. if the limit approaches infinity or negative infinity, then x=2 is your vertical asymptote...

OpenStudy (anonymous):

to find: x intercepts, plug in a 0 for y and solve for x; y intercepts, plug in a 0 for x and solve for y.

OpenStudy (anonymous):

@dpalnc, what about my graphing question?

OpenStudy (anonymous):

these question are all the things you need to create the graph...

OpenStudy (anonymous):

I know but how do I know whether to approach the HA from above or below

OpenStudy (anonymous):

first, do you know the limit of x/(x-2) as x approaches infinity...

OpenStudy (anonymous):

Well, as it approaches + infinity, The Limit is one, and as it approaches negative infinity it also is one

OpenStudy (anonymous):

ok so we know it only has 1 horizontal asymptote. now tell me... is there a solution to this equation: 1 = x/(x-2) ??

OpenStudy (anonymous):

look at the big picture... in order for that to be true, the numerator has to equal the denominator.. will that ever occur?

OpenStudy (anonymous):

No?

OpenStudy (anonymous):

correct... do you know what that means?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

it means our graph will not cross the horizontal asymptote (yes, horizontal asymptotes can be crossed). which means the graph will either be coming from the top and approach the asymptote or coming from the bottom and approach the asymptote...

OpenStudy (anonymous):

still with me?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

now, this is where our x/y intercepts come in.... what are they?

OpenStudy (anonymous):

x & y intercepts are both (0,0)

OpenStudy (anonymous):

ok.. what about our vertical asymptote(s)?

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

 this is what we have so far to construct the graph without the actual graph itself. but we actually have enough info to make a sketch of the function.. start with the x,y intercept...